Please answer the following question.

Please show all work.

A band pass communication link is similar to a base band system except that a carrier is introduced before transmission. The carrier is normally required because of A. antenna radiation requirements B. transmission line requirements. C. 1 Simulink simplifications. D. 1 data rate requirements
Answer #1

A. Antenna radiation requirements.

The primary reason for using high frequency carrier waves for sending information through EM waves is because it allows smaller antenna design. Typically, antenna length is given by L = wavelength/4

If carrier wave is not used, then for sending audio frequency of say 30 Hz, you will get quarter-wave antenna length of:
L = c/(4?f)=3?108/(4?30)=2.25?109metersc/(4?f)=3?108/(4?30)=2.25?109meters

A high frequency carrier would significantly reduce the antenna length.